How do I solve the system by substitution?
3x – y = –7
x + y = –9

Answer by checkley75(3666)

3X-Y=7
X+Y=-9
------------------------------------
4X=-2
X=2/-4
X=-1/2 ANSWER.
-1/2+Y=-9
Y=-9+1/2
Y=-17/2 ANSWER.
PROOF
-1/2-17/2=-9
-18/2=-9
-9=-9